OpenStudy (anonymous):

Learning about derivatives, instantaneous rate of change, ..via differential quotient equation.. and limits equations, I almost have it down now, need a review of the final few posts to check my work.. thanx Given: f[x] = x/e^x^2 f'[x] = Lim (h->0) [(f[x+h] -f[x])/ h] f'[x] = Lim (h->0) [ ((x+h)/(E^(x+h)^2) - (x/e^(x^2)) ) *1/h ]

OpenStudy (anonymous):

or at the very least please send me a good link that explains how to deal with canceling variables embedded in a polynomial exponent

jigglypuff314 (jigglypuff314):

Do you have to use the definition of the derivative to solve for your given function? derivative of e^u = d/du * e^u for example, derivative of e^(x^3) = (3x^2)*e^(x^3) Here's the proof for the concept if you would like: https://www.khanacademy.org/math/differential-calculus/taking-derivatives/der_common_functions/v/proof-d-dx-e-x-e-x

OpenStudy (anonymous):

\[f[x] = \frac{x}{e^{x^2}}\] \[ f'[x] = \lim_{h \rightarrow 0} \left( \frac{f[x+h] -f[x] }{h} \right) \] \[ f'[x] = \lim_{h \rightarrow 0} \left( \left( \frac{ (x+h) }{ E^{(x+h)^2} } - \frac{ x}{ E^{(x^2)} } \right) * \frac{1}{h} \right) \]

OpenStudy (anonymous):

Unfortunately I do..

OpenStudy (anonymous):

i have really come to admire those rules of the derivative..

OpenStudy (amistre64):

first principles eh

OpenStudy (anonymous):

yeah the baby steps..

OpenStudy (anonymous):

I've been on it for 3 days.. and my OCD wont allow me to let it go

OpenStudy (anonymous):

I cant friggin sleep, so pissed off.

OpenStudy (anonymous):

staring at equations for days. Im missing some fundamental cancelation trick of limits.. but all the vids Im watching, I still ahvent found one that deals with a complex rational like this, and a polynomial exponent.

OpenStudy (anonymous):

I can get there like this.. I know the answer to this is going to be \[ f'[x] = (1-2x^2)e^{x^2}\] and I can get there by.. g[x] = -x^2 g'[x] = -2x h[x] = E^g[x] h'[x] = -2xe^-x^2 f[x] = x h[x] f'[x] = (h'[x]/g'[x]) + (f[x] * g'[x]) but I cant submit this.

OpenStudy (anonymous):

and I cant reverse the logic, that would allow me to do that with algebra.

OpenStudy (australopithecus):

Its alright here, we can try to work through your final limit using this informal proof. \[\lim_{h \rightarrow 0} \frac{e^{-(x+h)^2}-e^{-x^2}}{h}\]

OpenStudy (australopithecus):

Although maybe amistre64 has a better way of solving this limit or problem

OpenStudy (anonymous):

Is that the first trick of limits, that we can remove x ? \[ f[x] = x/E^{x^2} \] \[ f[x] = x E^{x^2} \]

OpenStudy (anonymous):

oops .. that 's -x \[ f[x] = x E^{-x^2} \]

OpenStudy (australopithecus):

Well this problem demonstrates product rule as well, we already solved part of the derivative, we differentiated x/e^(x^2) in regards to the x term but we still need to differentiate x/e^(x^2) in terms of e^(-x^2) that is why we are left with that limit to solve

OpenStudy (australopithecus):

This is where you left off \[e^{-x^2}(xe^{x^2}(\lim_{h \rightarrow 0} \frac{e^{-(h+x)^2}- e^{-x^2}}{h}) + 1)\]

OpenStudy (anonymous):

oh, I just found your image...

OpenStudy (australopithecus):

yeah you could potentially use that to solve your problem

OpenStudy (australopithecus):

or the limit

OpenStudy (anonymous):

whats the strategy to get h out of that numerator ? should that even be our goal from here?

OpenStudy (australopithecus):

Well just considering the limit: \[\lim_{h \rightarrow 0} \frac{e^{-(h+x)^2}- e^{-x^2}}{h}\] Follow along with that image I posted you should be able to use it to solve this limit

OpenStudy (anonymous):

Whats the idea behind the first part of your image where f(x) becomes f(g(x)), is that because the equation x / e^x^2 has been split into two functions? And if so where was the split? was it at the x^2 ?

OpenStudy (australopithecus):

There is something you should note though that will likely come up the proof of d/dx e^x = e^x by definition of e: \[\lim_{h \rightarrow 0} \frac{e^h-1}{h} = 1\]

OpenStudy (australopithecus):

well e^(-x^2) is a composite function

OpenStudy (anonymous):

oh is that proof because technically, we have e^0 = (almost 1)-1 and we get a something/something

OpenStudy (australopithecus):

if f(x) = e^(x) and g(x) = -x^2 then f(g(x)) = e^(-x^2)

OpenStudy (australopithecus):

that is how the value of e is defined

OpenStudy (australopithecus):

one of the ways it is defined

OpenStudy (anonymous):

there was another x in there though.. f[x] = x / E^(x^2) is the x removed, or placed into f or g ?

OpenStudy (anonymous):

oops.. thats E^-x^2

OpenStudy (australopithecus):

we pulled it out of the limit

OpenStudy (australopithecus):

because it is technically a constant

OpenStudy (anonymous):

ah ok.. so that sits.. outside the lim definition still ? or does it now hold some value like 1?

OpenStudy (anonymous):

or does it cancel to 0 ? I havent clicked yet on how that constant rule is actually applied.

OpenStudy (anonymous):

just watching another vid on constant rule.

OpenStudy (anonymous):

lol, man, nobody explains this well.. they just gloss over and say ... 'oh we can pull this out.' and they move on

OpenStudy (anonymous):

okay, I think it goes outside ..

OpenStudy (australopithecus):

no it just sits outside the limit, if you have learned product rule this is a result of it. By product rule: d/dx f(x)g(x) = f'(x)g(x) + f(x)g'(x) in this problem assuming f(x) = e^(-x^2) and g(x) = x we solved f(x)g'(x) the limit left represents f'(x)g(x) since g(x) = x we can pull it out of the limit because it is just the original function, the limit itself represents f'(x)

OpenStudy (anonymous):

but it hangs around..

OpenStudy (australopithecus):

the e^(x^2) just cancels out, you could cancel it out right now it wouldn't change your solution

OpenStudy (anonymous):

so in reference to \[\lim_{h \rightarrow 0} \frac{e^h-1}{h} = 1\] To be clear, this means that if we encounter e^x then we have encountered the value 1? And anything attached to that or based on that can be considered to be factored with the value 1? so E^x^2 is actually 1^2 ?

OpenStudy (australopithecus):

No it is essentially what defines what the value of e is, it comes up when you try to use the definition of a derivative to find the derivative of e^x http://tutorial.math.lamar.edu/Classes/CalcI/DiffExpLogFcns.aspx

OpenStudy (anonymous):

hmmm, that's an odd and curious property. :/

OpenStudy (australopithecus):

I can help you work through the informal proof if you want.

OpenStudy (australopithecus):

@amistre64 , @Hero Might be able to help you more so with this.

OpenStudy (anonymous):

I think if the problem was just f[x] = e^(x^2) then I would have no problem.. x^2 = 2x and the derivative would be 2x * e^x^2 but I'm not sure what to do about that leading x/ .... maybe converting it into this form and then trying to work it out... f[x] = x e^-(x^2) do we just put the x outside the limit from the start? f'[x] = x Lim (h->0) [ (f[x+h] - f[x] ) /h ] It seems like its not a problem of 2 functions, but of 3 f[x] = x h[x] g[x] = E^h[x] h[x] = -x^2

OpenStudy (anonymous):

sorry thats x g[x]

OpenStudy (anonymous):

wish you could back edit these posts

OpenStudy (australopithecus):

Not sure what you are confused about here, look what I wrote about product rule. where g(x) = x and f(x) = e^(-x^2) the derivative of f(x)g'(x) is just e^(-x^2) This is shown in your problem, if you multiply e^(-x) by both terms left in the bracket you will get your x*(limit you need to solve) + e^(-x^2) the limit remaining is the derivative of f(x)

OpenStudy (australopithecus):

You would have this exact limit if you were trying to take the derivative with the definition using e^(-x^2)

OpenStudy (australopithecus):

If it was e^(x) and not e^(-x^2) you would have no problem because you wouldn't need chain rule

OpenStudy (australopithecus):

you are correct it is a problem of 3 functions

OpenStudy (australopithecus):

this problem requires product rule and chain rule

OpenStudy (australopithecus):

but the third function is easily differentiable so its not a problem

OpenStudy (australopithecus):

your problem is essentially m(x)*f(g(x))

OpenStudy (australopithecus):

this is actually a good thing for you to realize since you need to be able to recognize things like this when taking derivatives

OpenStudy (australopithecus):

you need to be able to split a function into multiple functions to take a derivative in most cases

OpenStudy (anonymous):

ah ok.. so I've been thinking g[x] = -x^2 for all cases of g[x] .. in this situation, and not seeing that g[x] is a term used to express generally a factor in an equation that varies from place to place.. where with Expression A (g[x]) * Expression B (f[x]), is in this case, x is in expression A and hence the derivative of g[x] * f[x] is in this case just E^-x^2 and in the next stage of this equation, g[x] can and will be defined to be something else, as it is applied to the general derivative equations.

OpenStudy (anonymous):

so Ive been scratching my head, trying to interpret the language

OpenStudy (australopithecus):

I am not understanding what you are saying. product rule is the following: f(x)*g(x) = f'(x)g(x) + f(x)g'(x) so if you had xln(x) the derivative would be: 1*ln(x) + (1/x)*x

OpenStudy (australopithecus):

g(x) = x g'(x) = 1 f(x) = ln(x) f'(x) = 1/x

OpenStudy (anonymous):

okay, I think I'm going to go read up more on the product rule and the chain rule again... by the way, is this problem going to need Ln[] to finish it off at some point?

OpenStudy (australopithecus):

I posted an informal proof you could follow along with

OpenStudy (australopithecus):

to solve the limit

OpenStudy (anonymous):

I was just saying before, that I wasn't understanding what and how g(x) was getting defined. And I just realized it was being defined on the fly, and changing in context of the problem as it evolved.

OpenStudy (anonymous):

so as you might understand from that last statement.. the informal proof had this mysterious g(x) throughout it.. and it wasn't intuitively clicking how that was defined and applied.

OpenStudy (australopithecus):

\[\lim_{h \rightarrow 0} \frac{e^{-(x + h)^2} - e^{-x^2}}{h} = \lim_{h \rightarrow 0} \frac{e^{-(x + h)^2} - e^{-x^2}}{h}*\frac{-(x+h)^2 - x^2}{-(x+h)^2 - x^2}\] do you understand that m(x) = e^(-x^2) is a composite function g(x) = -x^2, f(x) = e^x therefore, m(x) = f(g(x)) = e^(g(x)) = e^(-x^2)

OpenStudy (anonymous):

I think you've multiplied by 1, using the exponents to construct the 1.. you've chosen -x^2, so it looks like some kind of conjugate, constructed from the exponents.

OpenStudy (anonymous):

yes, I understand on the m(x) .. thnx

OpenStudy (australopithecus):

\[=(\lim_{h \rightarrow 0} \frac{e^{-(-x+h)^2}-e^{-x^2}}{-(x+h)^2 - x^2})*(\lim_{h \rightarrow 0} \frac{-(x+h)^2 - x^2}{h})\]

OpenStudy (australopithecus):

oops

OpenStudy (australopithecus):

I made a mistake

OpenStudy (australopithecus):

that should be +x^2

OpenStudy (australopithecus):

because - (-x^2) is + x^2

OpenStudy (australopithecus):

you should be able to solve the limit with that

OpenStudy (anonymous):

okay.. I'll have a go from there.. I got some visitors that just came in.. so give me an hour or so

OpenStudy (anonymous):

ok cool, they left.. I told em I was busy

OpenStudy (anonymous):

Are you saying this equation should look like this? \[ \lim_{h \rightarrow 0} \frac{e^{-(x + h)^2} - e^{-x^2}}{h}*\frac{-(x+h)^2 + x^2}{-(x+h)^2 + x^2} \]

OpenStudy (australopithecus):

Here \[(\lim_{h \rightarrow 0} \frac{e^{-(x+h)^2}-e^{-x^2}}{-(x+h)^2+ x^2})(\lim_{h \rightarrow 0} \frac{-(x+h)^2+ x^2}{h})\]

OpenStudy (anonymous):

ah because of the - in the top term .. cool

OpenStudy (australopithecus):

yeah then it splits into two limits

OpenStudy (australopithecus):

just solve them now they should be solvable

OpenStudy (australopithecus):

you will need to expand

OpenStudy (australopithecus):

everything

OpenStudy (anonymous):

that will be amazing

OpenStudy (anonymous):

Im excited

OpenStudy (anonymous):

\[ (\lim_{h \rightarrow 0} \frac{ e^{-(x+h)^2}-e^{-x^2} }{ -(x+h)^2+ x^2 } ) (\lim_{h \rightarrow 0} \frac{ -(x+h)^2+ x^2 }{ h } ) \] \[ (\lim_{h \rightarrow 0} \frac{ e^{-x^2-2hx-h^2}-e^{-x^2} }{ -x^2-2hx-h^2+ x^2 } ) (\lim_{h \rightarrow 0} \frac{ -x^2-2hx-h^2+ x^2 }{ h } ) \] so far ...

OpenStudy (australopithecus):

cancel stuff out

OpenStudy (anonymous):

\[(\lim_{h \rightarrow 0} \frac{ e^{-x^2-2hx-h^2}-e^{-x^2} }{ -2hx-h^2 } ) (\lim_{h \rightarrow 0} \frac{ -2hx-h^2 }{ h } )\] ... I think I'm going the wrong way with this.. should it be canceling within the exponent of the numerator somehow?

OpenStudy (anonymous):

this looks like it will zero out again. maybe if I multiply by log[] / log[] ? I dont know how else to get access to those exponents.. I can split them if they are E^(a + b) but E^(a - b) is a little trickier.

OpenStudy (anonymous):

I'll read on L hopitals. and the chain rule.. I cant help wondering if is requires some kind of Log[] / Log [] to pull that exponent out down out of the stratosphere. so we can get access to the terms.

OpenStudy (anonymous):

those negative exponents are hard to work with.. I know we can use a square root rule, for square roots, and Im not sure about higher degrees, but does that same idea apply to a polynomial in the exponent?

OpenStudy (anonymous):

maybe we can treat them like exponential functions in the global scope, and just throw away the low degree terms.

OpenStudy (anonymous):

or maybe we run the limit just on the exponent, and then after knowing which way the exponent goes.. we can then say what it will do as x when in e^(x)